OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (4, −2) and has a slope of −3? x + 3y = 10 x + 3y = −10 3x + y = 10 3x + y = −10

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ -2}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -3 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (anonymous):

x + 3y = 10 x + 3y = −10 3x + y = 10 3x + y = −10 ?? i dont get it

OpenStudy (anonymous):

i know it would be 3x because its the slope.

OpenStudy (jdoe0001):

right

OpenStudy (anonymous):

i dont see how 10 is the answer.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 4}}\quad ,&{\color{blue}{ -2}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -3 \\ \quad \\ y-{\color{blue}{ -2}}={\color{green}{ -3}}(x-{\color{red}{ 4}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \implies y+2=-3x+12\implies 3x+y=10\)

OpenStudy (anonymous):

thank you !

OpenStudy (jdoe0001):

I gather you weren't required to solve for "y", but anyway

OpenStudy (jdoe0001):

yw